Project

General

Profile

NDOS Commissioning Meeting 2011/04/28

Attendance: Jon, Tricia, Vince, Kurt, Enrique, Raphael, Carl, Mark Mathis, Andrew, Denis, Jim, Ron, Chad, Mat, Gavin, Minerba, Athans, Susan, Zeb, Jarek, Nick, Jan

Agenda:

  1. DCS EPICS presentation/demonstration
  2. DCS Plans
  3. DAQ/Electronics Update & Plans
  4. DataCheck update
  5. AOB

DCS EPICS presentation/demonstration

  • Jim K presented:
    • a couple of slides on what they have accomplished in the first 2 weeks; the presentation is in CVS
    • a demonstration of an actual working system on the teststand
  • Mark: if something goes out of tolerance, is the FEB protected by hardware? Andrew/Denis: yes
  • Andrew: what is the plan to roll this out to the DAQ cluster? How far away are you from being able to do this? Ron: code is already part of online suite, so there should not be much work to roll it out.
  • Jon: what are your next steps? Ron: needs to get IOC running during data collection? Jim: his involvement will rapidly decrease soon, Ron will continue to work on this.
    • error handling
    • database interface
    • integration with RMS messaging

DCS Plans

  • Chad: has been putting everything behind patch panels in racks 1 & 2. Things are looking nicer.
  • Athans:
    • posted doc-dB 5929, will present on Monday.
    • putting together a proposal for mechanical system readout (from Assembly group).
  • Denis: what water parameters do we measure in DCS? Chad: chiller temperature, pressure. Denis: we need flowmeters. Andrew: we should talk Mike Z. to find out what kind of flowmeters were installed.

DAQ/Electronics Update & Plans

  • Andrew: the TDU was moved to the MINOS service building today. This will add ~400 ns shift in time to data! Tomorrow we will hook up accelerator signals.

DataCheck Update: punt until Monday.

AOB